OpenStudy (anonymous):

simplify to standard form: -i (-4 + 2i) (6 - 7i) + 3 step by step

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know \(i^2=-1\)?

OpenStudy (anonymous):

misty! :D and noo ive forgotten a lot of things with math

OpenStudy (misty1212):

ok well it is true, \(i^2=-1\) lets take it one step at a time

OpenStudy (misty1212):

lets first compute \[-i(-4+2i)\]using the distributive law and see what we get as a first step it\[4i-2i^2\] right?

OpenStudy (misty1212):

this is where you say "yes, that is what you get when you distribute"

OpenStudy (anonymous):

lol!! I was looking at the equation to make sure and yes, that is what you get when you distribute

OpenStudy (misty1212):

ok and then we recall that \(i^2=-1\) so what is this really?

OpenStudy (anonymous):

my mind is in blank no lie :$

OpenStudy (misty1212):

what is \(-2i^2\) if \(i^2=-1\)

OpenStudy (anonymous):

I multiply -2 to -2 right?

OpenStudy (misty1212):

ahh nvm it is \(+2\) so you have \[(2+4i)(6-7i)+3\]

OpenStudy (anonymous):

okay so quick question so I wont forget and keep sounding dumb, whenever theres a lets say - i what can i replace it for?

OpenStudy (misty1212):

nothing

OpenStudy (misty1212):

\(-i=-i\) just leave it

OpenStudy (anonymous):

alright thanks I understood it a little bit more especially the positive 2 part

OpenStudy (misty1212):

your next job is \[(2+4i)(6-7i)\]

OpenStudy (anonymous):

6i -1i + 3? would it be something like that?